OpenStudy (anonymous):

In a large bag of candy, each of the five colors (red, orange, yellow, blue, and green) occurs with the same probability. You reach in and select 1. Find P (Red, yellow, or green). A) 0.04 B) 0.3 C) 0.6 D) 0.008

OpenStudy (campbell_st):

if there are 5 colours each with the same probability then P(any colour) = 1/5 or 0.2 P(Red, Yellow, Green) = P(red) + P(yellow)+P(green) = 0.2 + 0.2 + 0.2 just add the decimals for your answer